# Raman phonon dynamics and its control for enhanced optical frequency conversion

### 25 Mar 2026 v1

@@ 1 Introduction
Raman phonon represents the quantized motions in matter that originate from the inelastic Raman scattering of light [1, 2, 3, 4, 5, 6, 7]. The scattered light is redshifted through exchange of energy between photons and internal degrees of freedom of the medium, driving molecular motions, and thus inducing phonons. This Raman phonon excitation provides direct access to the microscopic dynamics of matter and forms the basis of a wide range of spectroscopic [8, 9, 10] and nonlinear optical phenomena [11, 12, 13].
@@
In addition to providing a fresh time-domain viewpoint, this framework reveals the linear phonon-mediated process embedded within Raman scattering. In a two-pulse configuration, phonons excited by the first pulse scatter the second pulse. Depending on the wave-vector matching between the pre-existing phonons and the Raman process driven by the second pulse, the scattering can produce Stokes [27, 28, 29, 30, 31] or anti-Stokes [32] radiation, corresponding to phonon generation or annihilation, respectively. Although this mechanism has been demonstrated experimentally, only recently have we provided a theoretical investigation [33, 14], and even then the discussion remained limited to two-pulse scenarios. Here, we further recognize that this linear phonon-mediated process constitutes the underlying origin of coherent Stokes and anti-Stokes Raman scattering (CSRS/CARS) [34, 35, 36, 37, 38, 39], whose well-known phase-matching condition originates from the phonon-mediated wave-vector matching. Moreover, this linear mechanism is distinct from the traditional interpretation based on nonlinear Raman (exponential) amplification. Due to the limitations of the traditional steady-state treatment, it has been historically overlooked. Notably, elucidating the underlying phonon interaction enables control of Raman generation beyond current capabilities. We introduce a phonon-controlled technique that suppresses early-stage cascaded Raman generation. By deliberately manipulating the light-phonon interaction, phonons generated by the first-order Stokes process exert minimal influence on subsequent Stokes generation, effectively confining the energy only to a specific order.
@@ 3 Linear phonon-mediated process
This process has long been considered as a four-wave-mixing (FWM) process, as in CSRS [34, 35, 36, 37] and CARS [38, 39]. In the traditional steady-state treatment, Raman temporal dynamics can be ignored and Raman-induced index modulation follows the pulse’s temporal profile, resembling the instantaneous electronic nonlinearity. In essence, the linear phonon-mediated process does involve four fields, two fields that produce the phonons and two fields in the latter phonon amplification or absorption process. Recent investigations in have demonstrated the importance of non-negligible temporal phonon dynamics due to the long-lived phonons [61, 62]. In particular, the probe pulse in CARS cannot be too strong; otherwise, its self-induced nonlinear Stokes generation can dominate over the linear anti-Stokes generation. This behavior is difficult to reconcile with a description based solely on the nonlinear FWM that also exponentially amplifies the field [as shown in Fig. 10(d) in [14]]. As a result, instantaneous FWM is an inaccurate description of this linear phonon-mediated process, and therefore also of the CSRS and CARS.
@@
This process also underlies the coherent Stokes generation from spontaneous Raman scattering in the transient regime, where the phonon lifetime is much longer than the pulse duration (Fig. 5) [71]. To investigate it, we simulated with a unidirectional pulse propagation (UPPE) that considers realistic Raman response function for several types of gases [14], and the result is shown in Fig. 5. Initially, the Stokes field is seeded by vacuum fluctuations [72]. Due to the incoherence of vacuum fluctuations, the beating with the pump field remains fluctuating, leading to incoherent phonon response. The Stokes increment during propagation [15] is therefore dominated by the pump field (top figure in Fig. 5). However, limited phonon response time (due to the Raman period) results in a smoothed phonon wave with increased coherence that further drives more coherent Stokes generation through the linear phonon-mediated process. This new coherent Stokes wave overrides the initial incoherent Stokes wave, seeding the latter nonlinear amplification and eventually producing a coherent Stokes field that inherits the temporal phase from the pump [Eq. (3a)]. Moreover, the growing phonon waves leads to a Stokes increment that is delayed to the trailing edge of the pump pulse. Traditional frameworks based on nonlinear Raman amplification, in which the incoherent Stokes background is amplified, cannot adequately account for the emergence of coherence in this process. et al.Carman analyzed transient Stokes generation by assuming a priori that the Stokes and pump fields share the same phase structure as they evolve in a high-gain environment [71]. Only with the introduction of the linear phonon-mediated dynamics is this coherent generation now clearly explained.
@@ 4 Phonon-controlled Raman scattering
To verify the idea of suppressing phonon-mediated processes, we simulated the vibrational Raman generation in -filled hollow-core fiber. Fiber and pulse parameters followed [41] that generates a frequency comb. A linearly-polarized input pulse was seeded by a weak Stokes pulse. Despite seeding, it undergoes significant early-stage thresholdless cascaded vibrational Raman scattering due to small wave-vector mismatch with first-order phonons across multiple Stokes orders (partially indicated by a small here) [Fig. 6(a)]. Stokes pulses not only experience reduced energy but also suffer degraded temporal quality, manifested as pronounced temporal substructures resulting from energy exchange among the various Stokes orders. In addition, we adjusted the gas pressure to yielding and thereby enabling efficient energy transfer from the first to the second Stokes order [Fig. 6(b)]. Unexpectedly, the increased wave-vector mismatch for the phonon-mediated second- to third-order process () suppresses further cascaded growth, causing Raman generation to terminate at the second Stokes order and resulting in highly-efficient second-order Stokes conversion.
@@
To assess its validity in the steady-state regime, we artificially reduced the Raman dephasing times while keeping the propagation distances fixed, and the resulting dynamics remain essentially unchanged (Fig. 7). This demonstrates that even in the steady-state limit where phonon temporal dynamics might appear negligible, the evolution is intrinsically transient: each temporal slice of the field is always influenced by phonons generated by earlier slices. More broadly, the universal action of the linear phonon-mediated process across all temporal regimes underscores that, irrespective of operating conditions, Raman scattering is fundamentally an ultrafast phonon-driven phenomenon governed by femtosecond- to picosecond-timescale phonon dynamics.

### 27 Mar 2026 v2

@@ 1 Introduction
Raman phonon represents the quantized motions in matter that originate from the inelastic Raman scattering of light [1, 2, 3, 4, 5, 6, 7]. The scattered light is redshifted through exchange of energy between photons and internal degrees of freedom of the medium, driving molecular motions, and thus inducing phonons. This Raman phonon excitation provides direct access to the microscopic dynamics of matter and forms the basis of a wide range of spectroscopic [8, 9] and nonlinear optical phenomena [10, 11, 12].
@@
In addition to providing a fresh time-domain viewpoint, this framework reveals the linear phonon-mediated process embedded within Raman scattering. In a two-pulse configuration, phonons excited by the first pulse scatter the second pulse. Depending on the wave-vector matching between the pre-existing phonons and the Raman process driven by the second pulse, the scattering can produce Stokes [26, 27, 28, 29, 30] or anti-Stokes [31] radiation, corresponding to phonon generation or annihilation, respectively. Although this mechanism has been demonstrated experimentally, only recently have we provided a theoretical investigation [32, 13], and even then the discussion remained limited to two-pulse scenarios. Here, we further recognize that this linear phonon-mediated process constitutes the underlying origin of coherent Stokes and anti-Stokes Raman scattering (CSRS/CARS) [33, 34, 35, 36, 37, 38, 39, 40], whose well-known phase-matching condition originates from the phonon-mediated wave-vector matching. Moreover, this linear mechanism is distinct from the traditional interpretation based on nonlinear Raman (exponential) amplification. Due to the limitations of the traditional steady-state treatment, it has been historically overlooked. Notably, elucidating the underlying phonon interaction enables control of Raman generation beyond current capabilities. We introduce a phonon-controlled technique that suppresses early-stage cascaded Raman generation. By deliberately manipulating the light-phonon interaction, phonons generated by the first-order Stokes process exert minimal influence on subsequent Stokes generation, effectively confining the energy only to a specific order.
@@ 3 Linear phonon-mediated process
This process has long been considered as a four-wave-mixing (FWM) process, as in CSRS [33, 34, 35, 36] and CARS [37, 38, 39, 40]. In the traditional steady-state treatment, Raman temporal dynamics can be ignored and Raman-induced index modulation follows the pulse’s temporal profile, resembling the instantaneous electronic nonlinearity. In essence, the linear phonon-mediated process does involve four fields, two fields that produce the phonons and two fields in the latter phonon amplification or absorption process. Recent investigations in have demonstrated the importance of non-negligible temporal phonon dynamics due to the long-lived phonons [62, 63]. In particular, the probe pulse in CARS cannot be too strong; otherwise, its self-induced nonlinear Stokes generation can dominate over the linear anti-Stokes generation [as shown in Fig. 10(d) in [13]]. This behavior is difficult to reconcile with a description based solely on the nonlinear FWM that also exponentially amplifies the field [64]. As a result, instantaneous FWM is an inaccurate description of this linear phonon-mediated process, and therefore also of the CSRS and CARS.
Raman gain suppression occurs when the wave-vector-matching condition is satisfied due to a delicate balance between phonon annihilation and creation, yielding zero nonlinear Raman gain [15, 65, 66, 67, 17]. It originates from the coupled action between instantaneous nonlinear phase modulation and Raman gain, and therefore cannot be neglected even in media with strong Raman nonlinearity, where the instantaneous contribution is dominated by the pulse-following Raman-induced index in the long-pulse regime [13]. In contrast, for linear phonon-mediated interactions, the same condition corresponds to simultaneous generation of both Stokes and anti-Stokes fields [Eq. (3)]. Pre-existing phonons, generated by a strong pump and its Stokes field, seed and sustain the scattering processes, overcoming the nonlinear Raman gain suppression. This is precisely the mechanism that enables degenerate CARS under perfect wave-vector matching.
@@
This process also underlies the coherent Stokes generation from spontaneous Raman scattering in the transient regime, where the phonon lifetime is much longer than the pulse duration (Fig. 5) [76]. To investigate it, we simulated with the unidirectional pulse propagation equation (UPPE) that considers realistic Raman response function for several types of gases [13], and the result is shown in Fig. 5. Initially, the Stokes field is seeded by vacuum fluctuations [77]. Due to the incoherence of vacuum fluctuations, the beating with the pump field remains fluctuating, leading to incoherent phonon response. The Stokes increment during propagation [14] is therefore dominated by the pump field (top figure in Fig. 5). However, limited phonon response time (due to the Raman period) results in a smoothed phonon wave with increased coherence that further drives more coherent Stokes generation through the linear phonon-mediated process. This new coherent Stokes wave overrides the initial incoherent Stokes wave, seeding the latter nonlinear amplification and eventually producing a coherent Stokes field that inherits the temporal phase from the pump [Eq. (3a)]. Moreover, the growing phonon wave leads to a Stokes increment that is delayed to the trailing edge of the pump pulse. Traditional frameworks based on nonlinear Raman amplification, in which the incoherent Stokes background is amplified, cannot adequately account for the emergence of coherence in this process. et al.Carman analyzed transient Stokes generation by assuming a priori that the Stokes and pump fields share the same phase structure as they evolve in a high-gain environment [76]. Only with the introduction of the linear phonon-mediated dynamics is this coherent generation now clearly explained.
Eq. (3) is introduced as an abstraction layer that isolates the linear phonon-mediated contribution from the full Raman response , clarifying the specific structural role this pathway plays in Raman temporal dynamics. Because the phonon wave is continuously driven by the optical fields, its amplitude and phase are not fixed material parameters but evolve with propagation distance. This construction therefore does not constitute an independent evolution equation; the actual dynamics must still be obtained from the full Raman formulation where explicitly contains both the field components that excite the phonons and those that are to be converted. The newly-introduced equations thus complement, rather than supplant, the conventional nonlinear description by selectively filtering out the linear phonon-mediated pathway to make its physical role more transparent.
@@ 4 Phonon-controlled Raman scattering
To verify the idea of suppressing phonon-mediated processes, we simulated the vibrational Raman generation in -filled hollow-core fiber with the aforementioned UPPE. Fiber and pulse parameters followed [42] that generates a frequency comb. A linearly-polarized input pulse was seeded by a weak Stokes pulse. Despite seeding, it undergoes significant early-stage thresholdless cascaded vibrational Raman scattering due to small wave-vector mismatch with first-order phonons across multiple Stokes orders (partially indicated by a small here) [Fig. 6(a)]. Stokes pulses not only experience reduced energy but also suffer degraded temporal quality, manifested as pronounced temporal substructures resulting from energy exchange among the various Stokes orders. In addition, we adjusted the gas pressure to yielding and thereby enabling efficient energy transfer from the first to the second Stokes order [Fig. 6(b)]. Unexpectedly, the increased wave-vector mismatch for the phonon-mediated second- to third-order process () suppresses further cascaded growth, causing Raman generation to terminate at the second Stokes order and resulting in highly-efficient second-order Stokes conversion.
@@
To assess its validity in the steady-state regime, we artificially reduced the Raman dephasing times while keeping the propagation distances fixed, and the resulting dynamics remain essentially unchanged (Fig. 7). This demonstrates that even in the steady-state limit where phonon temporal dynamics might appear negligible, the evolution is intrinsically transient: each temporal slice of the field is always influenced by phonons generated by earlier slices. More broadly, the universal action of the linear phonon-mediated process across all temporal regimes underscores that, irrespective of operating conditions, Raman scattering is fundamentally an ultrafast phonon-driven phenomenon governed by femtosecond phonon dynamics.